# Optimal ambition in business, politics and life

### v2

@@ Abstract
When should we be satisfied and when should we look for greener pastures? When is the perfect the enemy of the good? This type of question arises in many different contexts, including business, politics, resource exploitation, and our personal lives. Folk intuition suggests that people should aim for above-average results, but overreaching can lead to failure. Here, we mathematically formalize this intuition and relate it to empirical research across diverse domains. We model a search for strategies that have uncertain rewards over a fixed time period. The agent (i.e. searcher) knows the statistical distribution of rewards across strategies. At each time step, the agent either is satisfied and sticks with their current strategy or continues searching. We prove that the agent’s optimal satisfaction threshold is both finite and strictly larger than the mean of available rewards. Compared to the optimal threshold, being too ambitious has a higher expected cost than being too cautious, implying that uncertainty over the reward distribution should motivate caution. The optimal satisfaction threshold becomes larger if the search time is longer, or if the reward distribution is rugged (i.e., has low autocorrelation) or left skewed. Using upward social comparison to assess the reward landscape biases agents towards never being satisfied, which decreases their expected performance substantially. We discuss how these insights can be applied empirically, using examples from entrepreneurship, economic policy, political campaigns, online dating, and college admissions.
@@
Norman Vincent Peale famously said, “Shoot for the Moon. Even if you miss, you’ll land among the stars.” However, folk wisdom suggests that it is actually possible to aim too high, just as it is possible to aim too low. In a business context, entrepreneurs are encouraged to pursue big ideas, but not unrealistic pie-in-the-sky ones. Career advisers encourage people to build skills and social capital gradually, rather than expecting their first job to be perfect (e.g., Jay (2012)), and to be opportunistically mobile but not restless or non-committal. Common dating advice encourages people to have realistic expectations about what to look for in a partner. In politics, incrementalism encourages gradual changes to improve society without destabilizing its institutions Lindblom (1959). Dr. Martin Luther King Jr. famously said that “the arc of the moral universe is long, but it bends toward justice” King (March 31, 1968).
@@
Empirical research seems to support these folk intuitions encouraging intermediate ambition. For example, intermediate procrastinators are often the most successful entrepreneurs and innovators Grant (2017). They do not over-commit to their first ideas, but they still have the focus to execute better ideas later. Employees’ propensities to change jobs tend to have mixed effects on career success and professional well-being Guan et al. (2019). Research supports incrementalism as a policymaking strategy, due to checks and balances, distributed expertise, and pluralistic social and political interests Lindblom (1959); Hayes (2002). Utopianism has sometimes produced tragic consequences such as the Great Chinese Famine in Mao Zedong’s China Courtois (1999); Pinker (2012). The happiest marriages, on average, occur between people who are similar to each other in traits related to desirability (e.g., health, vitality, exercise, spirituality) George et al. (2015). Natural resource exploiters such as fishers, who do not have access to perfect information about where the resources are most plentiful, maximize their catch by aiming for patches of slightly above average quality Burgess et al. (2020a). These are a few examples of many.
@@
Here, we mathematically formalize this intuition. We show in a general search model that the optimal ambition targets outcomes that are both strictly finite and strictly above the mean of available rewards. In other words, the optimal ambition does not ‘shoot for the moon, but it does try to do better than average. We also show how optimal ambition depends on certain features of the reward landscape and how optimal ambition is affected by social comparison. In doing so, we provide a formal conceptual link between the diverse domains in which intermediate ambition seems to be supported, and we provide actionable insights into how to optimally calibrate ambition in these domains, using specific examples. Our model also provides testable hypotheses for future empirical research.
@@ I Model
We model an ‘agent’ searching among a set of possible strategies over a finite number of time periods, $t_{\text{max}}$. The agent does not know each strategy’s reward in advance, but they do have information about the statistical distribution of rewards across all strategies. We explore scenarios where the agents know the true distribution of rewards and where the agents estimate the distribution from their peers’ rewards.
Strategies and rewards could abstractly represent, for example, jobs and salaries for a job-seeker, colleges and application cycles for a prospective student, venture ideas and payoffs for an entrepreneur, dating prospects and relationship satisfaction, campaign strategies and popularity for a political candidate, policies and their outcomes for a policy maker, or patches with different harvests for a harvester. There are many other possible examples. Our model is admittedly simpler than these real decision settings, but we argue that its key insights should generalize to more complex settings.
During each time period, our agent either sticks with their current strategy or chooses a new one. Their objective is to maximize the sum of rewards across the $t_{\text{max}}$ time periods. Their key decision variable is when to be satisfied with the current strategy. We assume that agents have a satisfaction threshold, $T$, measured as a number of standard deviations, $\sigma$, above or below the mean reward across all strategies, $\mu$. The agent searches in each time step until they find a strategy with a reward greater than or equal to $T$. They stick with that strategy and its reward for the rest of the time periods. The higher $T$, the more ambitious the agent.
The agent must balance the costs of being too easily satisfied and settling for less than what is attainable, with the costs of being too ambitious and passing on high, attainable rewards while searching fruitlessly for the unattainable. We prove that the optimal satisfaction threshold, $T$, is above the mean, $\mu$ (agents should try to do better than average), but the optimal $T$ is also finite (overambition is costly). We explore how the optimal satisfaction threshold depends on several properties of the search, the reward distribution, and the information agents use to make their decisions.
@@
Here, $\epsilon_{t}$ are independent and identically distributed (i.i.d.) samples from the reward distribution. Our model is effectively an autoregressive (AR1) random walk, which ends at time $t_{\text{max}}$ ($t\leq t_{\text{max}}<\infty$). This approach has the advantage of model parsimony while capturing the properties of commonly used fitness landscapes (e.g., the underlying distribution Skellett et al. (2005) and parent–offspring correlation Verel et al. (2003) in fitness landscapes, and autocorrelation behavior on general neighborhood digraph landscapes Dimova et al. (2005)). (1) describes the two phases of the search. In the ‘explore’ phase, each agent conducts a random walk. Once the agent meets their respective threshold, they initiate the ‘exploit’ phase and remain at the same reward value until they run out of time.
@@ II Results
We can analytically derive the relationship between an agent’s satisfaction threshold and their expected reward, in the special case of a maximally rugged landscape ($\varphi=0$) and a Gaussian distribution of rewards, $\epsilon_{t}\sim\mathcal{N}(\mu=0,\sigma^{2})$ (see Materials and Methods). In other cases, we can accurately compute the relationship numerically (Fig. 2A).
@@
The expected reward has a unimodal (hump-shaped) relationship with the satisfaction threshold (Fig. 2A). Being always satisfied or never satisfied yields an expected cumulative reward of $\mu t_{\text{max}}$. Always-satisfied agents perpetually receive the reward of the first strategy they sample—the mean, $\mu$, on average. Never-satisfied agents randomly sample the distribution in each time period, earning an expected average reward of $\mu$. There is an intermediate range of satisfaction thresholds that earn higher expected rewards.
@@
Figure 2: A: analytical and numerical calculations of the expected reward distribution, as a function of the target threshold, on a maximally rugged landscape. The expected reward distribution is unimodal and negatively skewed, with an optimal threshold above zero (the landscape mean). The analytical expression given by (2) matches the simulation results. $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, $t_{\text{max}}=1000$, averaged over $10^{4}$ simulations. We note that many real-world applications occur on shorter time scales ($t_{\text{max}}\ll 1000$), and consequently have lower optimal thresholds (e.g., optimal $T=1.6$ with $t_{\text{max}}=100$). B and C: examples of real-world search strategies. B: When online dating, heterosexual men (blue) and women (red) are most likely to message potential partners who are slightly more desirable than they are (averaged over four cities shown in fig. 2 in ref. Bruch and Newman (2018)). C: When applying for college, members of the 2008 U.S. high-school graduating class concentrated their applications on schools with median SAT scores similar to, or slightly below, their own score (data from ref. Hoxby and Avery (2012)). This suggests students were either being sub-optimally ambitious, or faced other constraints (e.g., on income or geography).
@@
In the Materials and Methods section and Appendix A, we prove that the optimal threshold ($T$) is strictly greater than the mean reward ($\mu$). The proof is somewhat complex, but the intuition is simple. Suppose that there are two time steps ($t_{\text{max}}=2$). Searching in the second time step will yield an expected reward of $\mu$. Settling for anything less than $\mu$ after the first time step would therefore result in a lesser expected reward in the second time step.
@@
We also prove that the optimal threshold ($T$) increases in the search time ($t_{\text{max}}$), and is finite as long as the search time is finite (Materials and Methods and Appendix A). Again, both results are intuitive: more search time increases the likelihood and cumulative payoff of achieving an ambitious target.
@@
Overshooting the optimal threshold is costlier than undershooting. In other words, an agent overshooting the optimal threshold is expected to receive lower rewards than an agent undershooting the optimal threshold by the same amount (Figs. 2A, 3, and 4). This implies that uncertainty over the reward distribution reduces the optimal satisfaction threshold (i.e., motivates caution), similar to left-skewed uncertainty motivating risk aversion in other contexts Benuzzi and Ploner (2024).
@@
Fig. 2B and C illustrate two real-world contexts in which people seem to behave as though they understand the search tradeoff our theory captures. Fig. 2B compares the frequencies with which heterosexual men and women in four U.S. cities send messages to prospective partners on a dating app, as a function of their prospective partners’ desirability compared to their own (data from ref. Bruch and Newman (2018)). Long-term partners tend to have similar desirability George et al. (2015). Thus, a zero desirability gap might be analogous to the mean reward ($\mu$) in our model. In the online dating market, both men and women expend the greatest (modal) search effort on prospective partners slightly more desirable than themselves, as our theory would predict (Fig. 2B).
@@
Fig. 2C shows the distribution of college applications from the U.S. 2008 high-school graduating class, comparing students’ own scores on the Scholastic Aptitude Test (SAT) to those of the median current student at the college they were applying to (data from ref. Hoxby and Avery (2012)). Attending a college where one has the median SAT score might be analogous to the mean reward ($\mu$) in our model. As in the dating example, the modal search (application) effort occurred near this mean college. However, unlike the dating example, students sent more applications to weaker colleges than they sent to stronger ones. Hoxby Hoxby and Avery (2012) shows that most applications in this left tail came from low- and middle-income students, who faced greater non-academic constraints (e.g., geographic, economic) on which colleges they could attend. This low-ambition tail in college applications may also relate to the ‘safety school’ strategy, which prefers unambitious options in college to striking out and being left with no options.
@@ II.2 Reward landscape left skewness and ruggedness increase optimal satisfaction thresholds
Rugged landscapes create higher-variance rewards, which increases expected rewards and optimal satisfaction thresholds. Thus, agents should be more ambitious on rugged landscapes. The reward distribution for the first time period is the same for smooth and rugged landscapes, but autocorrelation reduces the subsequent variance on smooth landscapes (see Materials and Methods). This means that agents can only make incremental changes: setting their sights too high will likely result in never encountering the desired reward, which would require a gradual climb.
Holding the mean reward ($\mu$) constant, left (right) skew implies a higher (lower) median and mode reward (Fig. 1). Left-skewed landscapes therefore have higher optimal satisfaction thresholds, compared to the mean ($\mu$), since higher rewards are more abundant. Expected rewards are also higher in left-skewed reward landscapes, at the optimal threshold. Simulations in Fig. 3 assume skew-normal reward distributions, with mean of zero and variance of one, as in other simulations.
Figure 3: Smoother landscapes yield lower cumulative rewards than the maximally rugged, non-skewed landscapes (left). Left-skewed landscapes yield higher cumulative rewards (right). Smooth landscapes are generated via the analytical expression (see Methods). For the skewed landscape simulations, $\epsilon_{t}\sim\mathcal{SN}(0,1,\alpha)$, $\varphi=0$, $t_{\text{max}}=1000$, where the parameter $\alpha$ is varied to result in skew values from $-0.8$ to $0.8$, each averaged over $10^{4}$ simulations.

### v3

@@ Abstract
In business, politics and life, folk wisdom encourages people to aim for above-average results, but to not let the perfect be the enemy of the good. Here, we mathematically formalize and extend this folk wisdom. We model a time-limited search for strategies having uncertain rewards. At each time step, the searcher either is satisfied with their current reward or continues searching. We prove that the optimal satisfaction threshold is both finite and strictly larger than the mean of available rewards—matching the folk wisdom. We show that being too ambitious has a higher expected cost than being too cautious. We show that the optimal satisfaction threshold increases if the search time is longer, or if the reward distribution is rugged (i.e., has low autocorrelation) or left skewed. The skewness result reveals counterintuitive contrasts between optimal ambition and optimal risk taking. We show that using upward social comparison to assess the reward landscape substantially harms expected performance. We show how these insights can be applied in real-world settings, using examples from entrepreneurship, economic policy, political campaigns, online dating and college admissions.
@@
Norman Vincent Peale famously said, “Shoot for the Moon. Even if you miss, you’ll land among the stars.” However, folk wisdom suggests that it is actually possible to aim too high, just as it is possible to aim too low. In a business context, entrepreneurs are encouraged to pursue big ideas, but not unrealistic pie-in-the-sky ones. Career advisers encourage people to build skills and social capital gradually, rather than expecting their first job to be perfect (e.g., ref. Jay (2012)), and to be opportunistically mobile but not restless or non-committal. Common dating advice encourages people to have realistic expectations about what to look for in a partner. In politics, incrementalism encourages gradual changes to improve society without destabilizing its institutions Lindblom (1959). Dr. Martin Luther King Jr. famously said that “the arc of the moral universe is long, but it bends toward justice” King (March 31, 1968).
@@
Empirical research seems to support these folk intuitions encouraging intermediate ambition. For example, intermediate procrastinators are often the most successful entrepreneurs and innovators Grant (2017). They do not over-commit to their first ideas, but they still have the focus to execute better ideas later. Employees’ propensities to change jobs tend to have mixed effects on career success and professional well-being Guan et al. (2019). Research supports incrementalism as a policymaking strategy, due to checks and balances, distributed expertise, and pluralistic social and political interests Lindblom (1959); Hayes (2002). Utopianism has sometimes produced tragic consequences such as the Great Chinese Famine in Mao Zedong’s China Courtois (1999); Pinker (2012). The happiest marriages, on average, occur between people who are similar to each other in traits related to desirability (e.g., health, vitality, exercise, spirituality) George et al. (2015). Natural resource exploiters such as fishers, who do not have access to perfect information about where the resources are most plentiful, maximize their catch by aiming for patches of slightly above-average quality Burgess et al. (2020a). These are a few examples of many.
@@
Here, we mathematically formalize and extend these folk intuitions about ambition. We show in a general search model that optimal ambition targets outcomes that are both strictly finite and strictly larger than the mean of available rewards. In other words, optimal ambition does not ‘shoot for the moon’, but it does try to do better than average. Our model has conceptual parallels to areas of applied mathematics, economics, and theoretical biology related to optimal search and optimal stopping, which we discuss. We show how optimal ambition depends on certain features of the reward landscape, and how optimal ambition is affected by social comparison. Our results provide a precise and accessible conceptual link between the diverse domains in which intermediate ambition seems to be supported. Our model provides actionable insights for optimally calibrating ambition, which we illustrate using empirical examples from economic policy, entrepreneurship, and political campaigns. Our model also provides testable hypotheses for empirical research, which we illustrate using examples from online dating and college admissions.
@@ I Model
We model an ‘agent’ searching among a set of possible strategies over a finite number of time periods, $t_{\text{max}}$. The agent does not know each strategy’s reward in advance, but they do have information about the statistical distribution of rewards across all strategies. We explore scenarios where the agents know the true distribution of rewards, and scenarios where the agents estimate the distribution from their peers’ rewards.
Strategies and rewards could abstractly represent, for example, jobs and salaries for a job-seeker, colleges and application cycles for a prospective student, venture ideas and payoffs for an entrepreneur, dating prospects and relationship satisfaction for a single, campaign strategies and popularity for a political candidate, policies and their outcomes for a policy maker, or patches with different potential harvests for a harvester. There are many other possible examples. Our model is admittedly simpler than these real decision settings, but we argue that its key insights should generalize to more complex settings.
During each time period, our agent either sticks with their current strategy or chooses a new one. Their objective is to maximize the sum of rewards across the $t_{\text{max}}$ time periods. Their key decision variable is when to be satisfied with the current strategy. We assume that agents have a satisfaction threshold, $T$, measured as a number of standard deviations, $\sigma$, above or below the mean reward across all strategies, $\mu$. The agent searches in each time step until they find a strategy with a reward greater than or equal to $T$. Then, they stick with that strategy and its reward for the rest of the time periods. The higher $T$, the more ambitious the agent.
The agent must balance the costs of being too easily satisfied and settling for less than what is attainable, with the costs of being too ambitious and passing on high, attainable rewards while searching fruitlessly for the unattainable. We prove that the optimal satisfaction threshold, $T$, is strictly larger than the mean, $\mu$ (i.e., agents should try to do better than average), but the optimal $T$ is also finite (i.e., overambition is possible, and costly). We explore how the optimal satisfaction threshold depends on several properties of the search, the reward distribution, and the information agents use to make their decisions.
@@
Here, $\epsilon_{t}$ are independent and identically distributed (i.i.d.) samples from the reward distribution. Our model is effectively an autoregressive (AR1) random walk, which ends at time $t_{\text{max}}$ ($t\leq t_{\text{max}}<\infty$). This approach has the advantage of model parsimony while capturing the properties of commonly used fitness landscapes (e.g., the underlying distribution Skellett et al. (2005) and parent–offspring correlation Verel et al. (2003) in fitness landscapes, and autocorrelation behavior on general neighborhood digraph landscapes Dimova et al. (2005)). Equation (1) describes the two phases of the search. In the ‘explore’ phase, each agent conducts a random walk. Once the agent meets their respective threshold, they initiate the ‘exploit’ phase and remain at the same reward value until they run out of time.
Our model offers a general, tractable, and accessible formulation applicable to a widely studied class of sequential decision problems that aim to maximize rewards (or minimize costs) under uncertainty (e.g. MacQueen and Miller Jr (1960); van Moerbeke (1974); Freeman (1983); Powell (2019); Kohn and Shavell (1974)). Optimal stopping problems, concerned with determining the best time to take a particular action, have been widely applied in management science Ciocan and Mišić (2022) and finance Liu and Mu (2022). Optimal search has been explored in economics Kohn and Shavell (1974), and in theoretical biology via optimal foraging theory Pyke (1984). For example, the Marginal Value Theorem Charnov (1976) has been used to explain how long diving mammals spend underwater Kramer (1988), how long birds spend in any given tree Cowie (1977), and how long bumblebees spend on flowers Schmid-Hempel et al. (1985). These same ideas have even helped explain the evolution of adaptations Parker and Smith (1990). Depending on the context, solving such optimal stopping or optimal search problems may require advanced computational methods Becker et al. (2019); Ciocan and Mišić (2022). In contrast, in this work, we focus on a simple, self-contained formulation that emphasizes parsimony. For a detailed mathematical treatment of optimal stopping and hitting times in autoregressive processes, see refs.  Novikov (2009); Christensen (2012); Christensen et al. (2011). We extend our model to incorporate features that are psychologically relevant and commonly reflected in human decision-making, such as varying reward distribution propertiesGuan et al. (2014) and social comparison Guan et al. (2014); Gerber et al. (2018); Muller and Fayant (2010).
@@ II Results
We can analytically derive the relationship between an agent’s satisfaction threshold and their expected reward, in our model, in the special case of a maximally rugged landscape ($\varphi=0$) and a Gaussian distribution of rewards, $\epsilon_{t}\sim\mathcal{N}(\mu=0,\sigma^{2})$ (see Methods). In other cases, we can accurately compute the relationship numerically (Fig. 2a).
@@
The expected reward has a unimodal (i.e., hump-shaped) relationship with the satisfaction threshold (Fig. 2a). Being always satisfied or never satisfied yields an expected cumulative reward of $\mu t_{\text{max}}$. Always-satisfied agents perpetually receive the reward of the first strategy they sample—the mean, $\mu$, on average. Never-satisfied agents randomly sample the distribution in each time period, earning an expected average reward of $\mu$. There is an intermediate range of satisfaction thresholds that earn higher expected rewards.
@@
Figure 2: Panel a: analytical and numerical calculations of the expected reward distribution, as a function of the target threshold, on a maximally rugged landscape. The expected reward distribution is unimodal and negatively skewed, with an optimal threshold above zero (the landscape mean). The analytical expression given by Equation (2) (black, dashed) matches the simulation results (red). Parameter values are: $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, and $t_{\text{max}}=1000$. Results shown are averaged over $10^{4}$ simulations. We note that many real-world applications occur on shorter time scales ($t_{\text{max}}\ll 1000$), and consequently have lower optimal thresholds (e.g., optimal $T=1.6$ with $t_{\text{max}}=100$). Panels b and c: examples of real-world search strategies. Panel b: When online dating, heterosexual men (blue) and women (red) are most likely to message potential partners who are slightly more desirable than they are (averaged over the four cities shown in fig. 2 of ref. Bruch and Newman (2018)). Panel c: When applying for college, members of the 2008 U.S. high-school graduating class concentrated their applications on schools with median SAT scores similar to, or slightly below, their own score (data from ref. Hoxby and Avery (2012)). This suggests students were either being sub-optimally ambitious, or they faced other constraints (e.g., on income or geography), as ref. Hoxby and Avery (2012) suggests was the case.
@@
In the Methods section and Appendix A, we prove that the optimal threshold ($T$) is strictly greater than the mean reward ($\mu$). The proof is somewhat complex, but the intuition is simple. Suppose that there are two time steps ($t_{\text{max}}=2$). Searching in the second time step will yield an expected reward of $\mu$. Settling for anything less than $\mu$ after the first time step would therefore result in a lesser expected reward in the second time step.
@@
We also prove that the optimal threshold ($T$) increases in the search time ($t_{\text{max}}$), and is finite as long as the search time is finite (see Methods and Appendix A). Again, both results are intuitive: more search time increases the likelihood and cumulative payoff of achieving an ambitious target.
@@
Overshooting the optimal threshold is costlier than undershooting. In other words, an agent overshooting the optimal threshold is expected to receive lower rewards than an agent undershooting the optimal threshold by the same amount (Figs. 2a, 3, and 4). This implies that uncertainty over the reward distribution should reduce the optimal satisfaction threshold (i.e., it motivates caution), analogous to the way that left-skewed uncertainty motivates risk aversion in other contexts Benuzzi and Ploner (2024). However, in settings where agents learn about the reward distribution through searching and they can adjust their satisfaction thresholds accordingly, Kohn and Shavell Kohn and Shavell (1974) showed that uncertainty can increase the optimal satisfaction threshold due to the information value of searching.
@@
Fig. 2b and c illustrate two real-world contexts in which people seem to behave as though they understand the search tradeoff that our theory captures. Fig. 2b compares the frequencies with which heterosexual men and women in four U.S. cities sent messages to prospective partners on a dating app, as a function of their prospective partners’ desirability compared to their own desirability (data from ref. Bruch and Newman (2018)). Long-term partners tend to be similar in their desirability George et al. (2015). Thus, a zero desirability gap might be analogous to the mean reward ($\mu$) in our model. In the online dating market, both men and women expend the greatest (modal) search effort on prospective partners slightly more desirable than themselves, as our theory would predict (Fig. 2b).
@@
Fig. 2c shows the distribution of college applications from the 2008 U.S. high-school graduating class, comparing students’ own scores on the Scholastic Aptitude Test (SAT) to those of the median current student at the college they were applying to (data from ref. Hoxby and Avery (2012)). Attending a college where one has the median SAT score might be analogous to the mean reward ($\mu$) in our model. As in the dating example, the modal search (application) effort occurred near this mean college. However, unlike the dating example, students sent more applications to weaker colleges than they sent to stronger ones. Hoxby Hoxby and Avery (2012) showed that most applications in this left tail came from low- and middle-income students, who faced greater non-academic constraints (e.g., geographic, economic) on which colleges they could attend. This low-ambition tail in college applications may also relate to the ‘safety school’ strategy, which prefers unambitious options in college to striking out and being left with no options. College applicants are able to pursue multiple strategies (applications) simultaneously (albeit with some costs from time and application fees), unlike the agents in our model.
@@ II.2 Reward landscape left skewness and ruggedness increase optimal satisfaction thresholds
Rugged landscapes create higher-variance rewards, which increases achievable expected rewards and optimal satisfaction thresholds (Fig. 3a). Thus, agents should be more ambitious on rugged landscapes. The reward distribution for the first time period is the same for smooth and rugged landscapes, but autocorrelation reduces the subsequent variance on smooth landscapes (see Methods). This means that agents can only make incremental changes: setting their sights too high will likely result in never encountering the desired reward, which would require a gradual climb due to the aurocorrelation.
Holding the mean reward ($\mu$) constant, left (right) skew implies a higher (lower) median and mode reward (Fig. 1). Left-skewed reward landscapes therefore have higher optimal satisfaction thresholds, compared to the mean ($\mu$), since higher rewards are more abundant (Fig. 3b). Expected rewards are also higher in left-skewed reward landscapes, at the optimal threshold. Right-skewed reward landscapes have the opposite properties. Simulations in Fig. 3b assume skew-normal reward distributions, with mean of zero and variance of one, as in other simulations.
Figure 3: Smoother reward landscapes yield lower cumulative rewards than the maximally rugged, non-skewed landscapes (a). Left-skewed reward landscapes yield higher cumulative rewards (b). Smooth landscapes are generated analytically (see Methods). For the skewed landscape simulations, $\epsilon_{t}\sim\mathcal{SN}(0,1,\alpha)$, $\varphi=0$, $t_{\text{max}}=1000$, where the parameter $\alpha$ is varied to result in skew values from $-0.8$ to $0.8$, each averaged over $10^{4}$ simulations.